# Regional chemical potential analysis for material surfaces

### v1

@@ III Simple model
$\psi(r)$ $=\sqrt{n(r_{s})}e^{-\alpha(r-r_{s})}=Ae^{-\alpha r},$ (16)
where $A=\sqrt{n(r_{s})}e^{\alpha r_{s}}$, $r_{s}$ is the distance between the material surface and the nucleus position $r=0$. The kinetic energy density and the electronic stress energy density are represented as
$T(r)$ $=-\frac{\hbar^{2}}{2m}\left(\alpha^{2}-\alpha\frac{2}{r}\right)A^{2}e^{-2\alpha r},$ (17)
@@
The RCP at the material surface is given by
@@
The above linear response approximation is reasonable, since we assumed that this system has only one electron. From Eq. (19), we can find $\mu_{R}^{\tau}$ at the surface is directly related with the decay constant $\alpha$ of the $s$-type STO. It can be a measure of delocalization of electrons from near the surface outward. By calculating the values of $r_{s}$ and $\mu_{R}^{\tau}(r_{s})$ from a DFT result, the value of $\alpha$ can be estimated from Eq. (19). For the hydrogen atom, the DFT result gives $r_{s}=2.0$[a.u.], $\mu_{R}^{\tau}(r_{s})=-0.23$[a.u.] and $\alpha=0.92$. The computational details of DFT calculations will be explained in a later section.
As a next step, we assume a H2 molecular model to investigate the chemical bonding force between two atomic models separated by the distance $l$. Using a linear combination of the spherical STOs, the H2 molecular model can be written as
@@
$\psi_{\rm s}$ $=\sqrt{n_{s}}e^{-\alpha_{s}(z-z_{s})}\Lambda_{s}(r,\theta),$ (23)
$\psi_{\rm t}$ $=\sqrt{n_{t}}e^{\beta_{t}(z-z_{t})}\Lambda_{t}(r,\theta),$ (24)
where $n_{s}$ and $n_{t}$ are the electron number density at the material surface ($z=z_{s},z_{t}$), which is defined by the zero kinetic energy density of the single atomic model. This model is considered to be based on the assumption that the atom $s$ on $z=0$ represents an atom of a material surface and the atom $t$ on $z=l$ represents an apex atom of a tip used in AFM experiments as shown in Fig. 2 (a). $\Lambda_{s}(r,\theta)$ and $\Lambda_{s}(r,\theta)$ are functions of the polar coordinates. Later, we will see that this simpler model is enough to reproduce a force curve of H2 molecule.
@@
$E_{\pm}\approx\frac{1}{2}\Bigg{(}({\cal E}_{s}+{\cal E}_{t})\pm\sqrt{({\cal E}_{s}-{\cal E}_{t})^{2}+\left(-\left({\cal E}_{s}-{\cal E}_{t}\right)^{2}+({\cal E}_{s}+{\cal E}_{t})^{2}(K-1)^{2}\right)\tilde{S}}\Bigg{)}.$ (28)
The plus (minus) sign corresponds to the solution of covalent (anti covalent) bonding. Furthermore, we assume that ${\cal E}_{s}={\cal E}_{t}$, $\alpha_{s}=\beta_{t}$, and $n_{s}=n_{t}$. The dependency of the distance $l$ is included in $\tilde{S}$ through the overlap matrix $S_{st}$. Therefore, the two eigen energies $E_{\pm}$ are approximately given as
@@
where $F(l)$ represents a force between two atoms separated by the distance $l$. In the case of the spherical STO model $\Psi^{\rm sph}$, substituting $\tilde{S}=\left(1+\alpha_{s}l+\frac{1}{3}(\alpha_{s}l)^{2}\right)^{2}e^{-2\alpha_{s}l}$ into Eq. (31),
@@
where the $r_{\rm cut}$ is a parameter for the cutoff radius. The Eqs. (32) and  (33) reveal that the shape of the force curve is determined by the value of $\alpha_{s}$. Since $\mu_{R}^{\tau}(r_{s})$ gives $\alpha_{s}$ via Eq. (19), the $\mu_{R}^{\tau}(\vec{r})$ at material surfaces can be a measure of the strength of the chemical bonding force. This fact supports that a higher value of $\mu_{R}^{\tau}(\vec{r})$ at the surface suggests a stronger electron-donating ability toward the bonding region, which contributes to covalent bonding, as discussed in the previous section.
A comparison of the above Eqs. (32) and  (33) with a DFT result is shown in Fig. 2 (b). The parameters, ${\cal E}_{s}\approx\mu_{R}^{\tau}(r_{s})=-0.23$[a.u.], $\alpha_{s}=0.92$, $K=1.75$ [23], and $\sqrt{\frac{|\Lambda_{st}|^{2}}{\Lambda_{ss}\Lambda_{tt}}}\approx 1$ were used for plotting the Fig. 2(b). The values of $\mu_{R}^{\tau}(r_{s})$ and $\alpha_{s}$ were obtained by the hydrogen atom model. Considering that the atomic surfaces are almost in contact at $l=2r_{s}=2.11$, these models are in very good agreement at the chemical bonding region, $l>2.11$.
In this section, we revealed the relation between $\mu_{R}^{\tau}(r_{s})$ and the force $F(l)$ using a simple H2 model. However, the approximation in Eq. (19) cannot be applied to general materials. The extension of the approximation of $\mu_{R}^{\tau}(\vec{r})$ in general materials is devoted to the next section.
Figure 2: (a) A diatomic molecular model and its wave functions. (b) Force curves of DFT (green), spherical STO model (blue) and cylindrical STO model (red). The parameters, ${\cal E}_{s}\approx\mu_{\tau}(r_{s})=-0.23$[a.u.], $\alpha_{s}=0.92$, $K=1.75$ [23], and $\sqrt{\frac{|\Lambda_{st}|^{2}}{\Lambda_{ss}\Lambda_{tt}}}\approx 1$ were used.
@@ IV Approximation for regional chemical potential
$n_{\rm ew}(\vec{r})$ $=\int_{-\infty}^{\infty}dEn(\vec{r},E)\left[f(E,\mu_{\rm upper},\tilde{T}_{\rm upper})-f(E,\mu_{\rm lower},\tilde{T}_{\rm lower}))\right],$ (36)
$\varepsilon_{\tau,\rm ew}(\vec{r})$ $=\int_{-\infty}^{\infty}dE\varepsilon_{\tau}(\vec{r},E)\left[f(E,\mu_{\rm upper},\tilde{T}_{\rm upper})-f(E,\mu_{\rm lower},\tilde{T}_{\rm lower}))\right],$ (37)
where the Fermi function is defined as $f(E,\mu,\tilde{T})\equiv 1/(1+e^{(E-\mu)/\tilde{T}})$ as shown in Fig. 3. The energy dependent electron number density $n(\vec{r},E)$ and regional energy density $\varepsilon_{\tau}(\vec{r},E)$ are given in Appendix A. The four parameters $\mu_{\rm lower}$, $\mu_{\rm upper}$, $\tilde{T}_{\rm lower}$, and $\tilde{T}_{\rm upper}$ should be selected properly for the purpose of physical considerations.
Figure 3: An example of the Fermi functions $f_{\rm upper}=f(E,\mu_{\rm upper},\tilde{T}_{\rm upper})$ and $f_{\rm lower}=f(E,\mu_{\rm lower},\tilde{T}_{\rm lower})$, where $\mu_{\rm lower}=0$, $\mu_{\rm upper}=5$, $\tilde{T}_{\rm lower}=\tilde{T}_{\rm upper}=1$.
@@ V Computational details
The first-principles calculations based on DFT were carried out for H atom, H2 molecule, some organic molecules, and C(001) diamond surface. The DFT calculations within a generalized gradient approximation (GGA) [24, 25] were performed for the geometry relaxations and variable cell optimizations using the OpenMX code [26] based on norm-conserving pseudopotentials generated with multireference energies [27] and optimized pseudoatomic basis functions [28]. The standard basis set we used are listed in OpenMX website [29]. The qualities of the basis functions and fully relativistic pseudopotentials were carefully benchmarked by the delta gauge method [30] to ensure the accuracy of our calculations. An electronic temperature of 300 K is used to count the number of electrons by the Fermi-Dirac function. The regular mesh of 240 Ry in real space was used for the numerical integration and for the solution of the Poisson equation [31, 32]. Cell vectors and internal coordinates are simultaneously optimized without any constraint by using a combination scheme of the rational function (RF) method [33] and the direct inversion iterative sub-space (DIIS) method [34] with a BFGS update [35, 36, 37, 38] for the approximate Hessian. The force on each atom was relaxed to be less than 0.0003 Hartree/Bohr. For the C(001) diamond surface calculations, we used the same computational conditions in Ref. [20].
Calculations of the local RCPs and the kinetic energy density were performed by FLPQ module [39] in QEDalpha package [40]. FLPQViewer was used for visualization of the atomic structures, RCPs, and other properties. Similar functions are available in 3D Viewers such as OpenMX Viewer [41] and Fermisurfer [42]. The atomic unit is used for $\mu_{R}^{\tau}(\vec{r})$ (i.e. 1[Hartree/Bohr3] = 183.631[eV/Å3]). The broadening factors appeared in Eqs. (36) and (37) were set to $\tilde{T}_{\rm lower}=\tilde{T}_{\rm upper}=0.001/k_{B}[{\rm K}]$ for all the systems, where the Boltzmann constant $k_{B}=8.617\times 10^{-5}[{\rm eV}/{\rm K}]$.
@@ VI.1 C2H6, C2H4, and C2H5 molecules
Figures 4 (a)-(c) show the energy levels of C2H6, C2H4, and C2H5 molecular models, respectively. To draw figures of each local RCP on the surfaces $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$, at least HOMO (Highest Occupied Molecular Orbital) should be included in the energy window $\mu_{\rm lower}<\mu<\mu_{\rm upper}$. Therefore, the energy windows were chosen to be -5 eV to 0 eV for C2H6, -3 eV to 0 eV for C2H4, and -2 eV to 0 eV for C2H5. The corresponding $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ are shown in Figs. 4 (d)-(f).
In Fig. 4 (d), though the H atoms are visualized as green parts, the overall $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ is small. On the other hand, the C2H4 (Fig. 4 (e)) has the high $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ (red regions) originated from a C-C double bond. This result is consistent with the fact that the double bonding region is a highly electron-donating region, where electrons are more likely to be supplied to outer regions. Figure 4 (f) shows a molecule model whose structure is generated as one H atom is deleted from C2H6 without geometry optimization. The red regions at right top in the C2H5 correspond to the high $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ that originates from a dangling bond, which was bonded with a deleted H atom. This fact is consistent with the HOMO distribution depicted in Fig. 4 (f).
Figure 4: Energy levels of (a) C2H6, (b) C2H4, and (c) C2H5 molecular models. The area colored in orange represents the energy window for local RCP $\mu_{R}^{\tau}(\vec{r})$. The RCP on the surface for (d) C2H6, (e) C2H4, and (f) C2H5 molecular models are drawn as color maps on the surface. The HOMO of C2H5 are drawn as the red (+0.05 a.u.) and blue (-0.05 a.u.) isosurfaces.
VI.2 Benzene, naphtalene, and pentacene molecules
The RCP $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ of benzene (C6H6), naphthalene (C10H8) and pentacene (C22H14) molecules are shown in Fig. 5 (d)-(g), and the corresponding energy window is shown in Figs. 5 (a)-(c). For pentacene, RCP $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ in the energy range between -2 eV and 0 eV is shown in Fig. 5 (e), and one in the energy range between 0 eV and 2 eV is shown in Fig. 5 (f), separately.
Benzene is known to have delocalized electrons due to the $\pi$ bonds created by the six-membered ring. Correspondingly, $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ is high throughout the benzene ring as shown in Fig. 5 (d). On the other hand, in the case of naphtalene shown in Fig. 5 (e), the high $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ due to the double bond is seen at four corners at the upper and lower ends. As easily expected, the RCP $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ originating from the Lowest Unoccupied Molecular Orbital (LUMO) contribution shown in Fig. 5 (f) is totally different from that originating from the HOMO contribution shown in Fig. 5 (e). Therefore, it is necessary to determine the appropriate energy window depending on whether the target substance is a donor or acceptor of electrons. It is clear from the $\mu_{R}^{\tau}(\vec{r})$ distribution shown in Fig. 5 (g) that pentacene has a bonding state between the outer central carbons similar to that of benzene rings, and four double bonding sites similar to those of naphthalene at the upper and lower ends, which is also consistent with the bond length argument. When molecules interact with another inactive material surface such as an AFM probe with CO molecules attached, high $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ becomes a measure of the Pauli repulsive force originating from antibonding. The Fig. 5 (g) is consistent with the topology image of non-contact AFM experiments using a CO probe [21, 43].
@@ VI.2 Benzene, naphtalene, and pentacene molecules
VI.3 Benzene substituents
The energy levels, Hartree potentials at material surfaces, and the RCP $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ of phenol (C6H5OH) and nitrobenzene (C6H5NO2) molecules are shown in Fig. 6. It is known that substituted benzenes exhibit distinct regioselectivities in electrophilic aromatic substitution reactions. Phenol acts as an ortho/para-directing group due to the electron-donating effect of the hydroxyl substituent, whereas nitrobenzene is meta-directing as a result of the electron-withdrawing nature of the nitro group [44]. The electrostatic potential (Hartee potential), which is caused by intramolecular polarization, is a measure of Coulomb interaction with charged particles. Although the Hartree potential is sometimes used for discussion of electron-donating ability for ionic materials, it is not effective for regioselectivities in selectrophilic aromatic substitution reactions as shown in Figs. 6 (c) and (d). On the other hand, the local RCPs indicate the electron-donating ability at the surface. The distribution of local RCPs clearly visualizes the ortho/para-directing of the phenol and the meta-directing of the nitrobenzene. Here, we note that the energy windows were selected to include multiple orbitals, as it is known that the HOMO alone cannot account for the meta-directing behavior of nitrobenzene [45].

### v2

@@ III Simple model
$\psi^{\rm STO}(r)$ $=\sqrt{n(r_{s})}e^{-\alpha(r-r_{s})}=Ae^{-\alpha r},$ (17)
where $A=\sqrt{n(r_{s})}e^{\alpha r_{s}}$. The parameter $r_{s}$ is the distance between the nucleus position ($r=0$) and the electronic interface, which is the material surface defined by the isosurface of the zero kinetic energy density as discussed in Sec. II.2. Substituting $\psi^{L}=\left(\begin{smallmatrix}\psi^{\rm STO}(r)\\ 0\end{smallmatrix}\right)$ into Eqs. (9), (14), and (15), the kinetic energy density $T_{e}(r)$ and the regional energy density $\varepsilon_{\tau}(r)$ are represented as
$T_{e}(r)$ $=-\frac{\hbar^{2}}{2m}\left(\alpha^{2}-\alpha\frac{2}{r}\right)A^{2}e^{-2\alpha r},$ (18)
@@
The RCP at the electronic interface is given by
@@
The above linear response approximation is reasonable, since we assumed that this system has only one electron. From Eq. (20), we can find $\mu_{R}^{\tau}$ at the surface is directly related to the decay constant $\alpha$ of the $s$-type STO. It can be a measure of delocalization of electrons from near the surface outward. By calculating the values of $r_{s}$ and $\mu_{R}^{\tau}(r_{s})$ from a DFT result of a hydrogen atom, the value of $\alpha$ can be estimated from Eq. (20). For the hydrogen atom, the DFT result gives $r_{s}=2.0$[a.u.], $\mu_{R}^{\tau}(r_{s})=-0.23$[a.u.] and $\alpha=0.92$. This value of $\alpha$ is subsequently applied in the force calculation for an H2 molecular model described in the next paragraph. The computational details of DFT calculations will be explained in a later section.
As a next step, we assume an H2 molecular model to investigate the chemical bonding force between two atomic models separated by distance $l$. When the $\psi^{\rm STO}$ of each atom is constructed from DFT calculations of an isolated hydrogen atom, using a linear combination of the spherical STOs, the H2 molecular model can be written as
@@
Figure 2: (a) A diatomic molecular model and its wave functions. (b) Force curves of DFT (green), spherical STO model (blue) and cylindrical STO model (red). The parameters, ${\cal E}_{s}\approx\mu_{R}^{\tau}(r_{s})=-0.23$[a.u.], $\alpha_{s}=0.92$, $K=1.75$ [24], and $\sqrt{\frac{|\Lambda_{st}|^{2}}{\Lambda_{ss}\Lambda_{tt}}}\approx 1$ were used.
@@
$\psi_{s}$ $=\sqrt{n_{s}}e^{-\alpha_{s}(z-z_{s})}\Lambda_{s}(r,\theta),$ (24)
$\psi_{t}$ $=\sqrt{n_{t}}e^{\beta_{t}(z-z_{t})}\Lambda_{t}(r,\theta),$ (25)
where $n_{s}$ and $n_{t}$ are the electron number density at the electronic interface ($z=z_{s},z_{t}$), which is defined by the zero kinetic energy density of the single atomic model. This model is considered to be based on the assumption that the atom $s$ on $z=0$ represents an atom of a material surface and the atom $t$ on $z=l$ represents an apex atom of a tip used in AFM experiments as shown in Fig. 2 (a). $\Lambda_{s}(r,\theta)$ and $\Lambda_{t}(r,\theta)$ are functions of the polar coordinates. Later, we will see that this simpler model is enough to reproduce a force curve of H2 molecule.
@@
$E_{\pm}\approx\frac{1}{2}\left({\cal E}_{s+t}\pm\sqrt{{\cal E}_{s-t}^{2}+\left(-{\cal E}_{s-t}^{2}+{\cal E}_{s+t}^{2}(K-1)^{2}\right)\tilde{S}}\right),$ (29)
where ${\cal E}_{s\pm t}\equiv{\cal E}_{s}\pm{\cal E}_{t}$. The negative (positive) sign in front of the square root corresponds to the solution of covalent (anti-covalent) bonding. Furthermore, we assume that ${\cal E}_{s}={\cal E}_{t}$, $\alpha_{s}=\beta_{t}$, and $n_{s}=n_{t}$. The dependency of the distance $l$ is included in $\tilde{S}$ through the overlap matrix $S_{st}$. Therefore, the two eigenenergies $E_{\pm}$ are approximately given as
@@
where $F(l)$ represents a force between two atoms separated by the distance $l$.
In the case of the spherical STO model $\Psi^{\rm sph}$, substituting $\tilde{S}=\left(1+\alpha_{s}l+\frac{1}{3}(\alpha_{s}l)^{2}\right)^{2}e^{-2\alpha_{s}l}$ into Eq. (32),
@@
where $r_{\rm cut}$ is a parameter for the cutoff radius. Eqs. (33) and (34) reveal that the shape of the force curve is determined by the value of $\alpha_{s}$. Since $\mu_{R}^{\tau}(r_{s})$ gives $\alpha_{s}$ via Eq. (20), $\mu_{R}^{\tau}(\vec{r})$ at electronic interfaces can be a measure of the strength of the chemical bonding force. This fact supports that a higher value of $\mu_{R}^{\tau}(\vec{r})$ at the surface suggests a stronger electron-donating ability toward the bonding region, which contributes to covalent bonding, as discussed in the previous section.
A comparison of the above Eqs. (33) and (34) with a DFT result is shown in Fig. 2 (b). The parameters, ${\cal E}_{s}\approx\mu_{R}^{\tau}(r_{s})=-0.23$[a.u.], $\alpha_{s}=0.92$, $K=1.75$ [24], and $\sqrt{\frac{|\Lambda_{st}|^{2}}{\Lambda_{ss}\Lambda_{tt}}}\approx 1$ were used for plotting Fig. 2(b). The values of $\mu_{R}^{\tau}(r_{s})$ and $\alpha_{s}$ were obtained by the hydrogen atom model. Considering that the atomic surfaces are almost in contact at $l=2r_{s}=2.11$, these models are in very good agreement at the chemical bonding region, $l>2.11$.
In this section, we have revealed the relation between $\mu_{R}^{\tau}(r_{s})$ and the force $F(l)$ using a simple H2 model. However, the approximation in Eq. (20) cannot be applied to general materials. The extension of the approximation of $\mu_{R}^{\tau}(\vec{r})$ in general materials is discussed in the next section.
@@ IV Approximation for regional chemical potential
[…]
where the upper and lower Fermi functions are defined as $f_{\lambda}(E)\equiv f(E,\mu_{\lambda},\tilde{T}_{\lambda})\equiv 1/(1+e^{(E-\mu_{\lambda})/\tilde{T}_{\lambda}})$ as shown in Fig. 3. The energy dependent electron number density $n(\vec{r},E)$ and regional energy density $\varepsilon_{\tau}(\vec{r},E)$ are given in Appendix A. The four parameters $\mu_{\rm lower}$, $\mu_{\rm upper}$, $\tilde{T}_{\rm lower}$, and $\tilde{T}_{\rm upper}$ should be properly selected for the purpose of physical considerations.
Figure 3: An example of the Fermi functions $f_{\rm upper}$ and $f_{\rm lower}$. The parameters are set as $\mu_{\rm lower}=0$, $\mu_{\rm upper}=5$, $\tilde{T}_{\rm lower}=\tilde{T}_{\rm upper}=1$.
Figure 4: Energy levels of (a) C2H6, (b) C2H4, and (c) C2H5 molecular models. The area colored in orange represents the energy window for local RCP $\mu_{R}^{\tau}(\vec{r})$. The RCP on the surface for (d) C2H6, (e) C2H4, and (f) C2H5 molecular models are drawn as color maps on the surface. (g) The HOMO of C2H5 are drawn as the red (+0.05 a.u.) and blue (-0.05 a.u.) isosurfaces.
@@ V Computational details
The first-principles calculations based on DFT were carried out for H atom, H2 molecule, some organic molecules, and C(001) diamond surface. The DFT calculations within a generalized gradient approximation (GGA) [25, 26] were performed for the geometry relaxations and variable cell optimizations using the OpenMX code [27] based on norm-conserving pseudopotentials generated with multireference energies [28] and optimized pseudoatomic basis functions [29]. The standard basis set we used are listed in OpenMX website [30]. The qualities of the basis functions and fully relativistic pseudopotentials were carefully benchmarked by the delta gauge method [31] to ensure the accuracy of our calculations. An electronic temperature of 300 K is used to count the number of electrons by the Fermi-Dirac function. The regular mesh of 240 Ry in real space was used for the numerical integration and for the solution of the Poisson equation [32, 33]. Cell vectors and internal coordinates are optimized by using a combination scheme of the rational function (RF) method [34] and the direct inversion iterative sub-space (DIIS) method [35] with a BFGS update [36, 37, 38, 39] for the approximate Hessian. The force on each atom was relaxed to be less than 0.0003 Hartree/Bohr. For the C(001) diamond surface calculations, we used the same computational conditions in Ref. [21].
Calculations of the local RCPs and the kinetic energy density were performed by FLPQ module [40] in QEDalpha package [41]. FLPQViewer [42] was used for visualization of the atomic structures, RCPs, and other properties. Similar functions are available in 3D Viewers such as OpenMX Viewer [43] and Fermisurfer [44]. The atomic unit is used for $\mu_{R}^{\tau}(\vec{r})$ (i.e. 1[Hartree/Bohr3] = 183.631[eV/Å3]). The broadening factors appeared in Eqs. (37) and (38) were set to $\tilde{T}_{\rm lower}=\tilde{T}_{\rm upper}=0.001/k_{B}[{\rm K}]$ for all the systems, where the Boltzmann constant $k_{B}=8.617\times 10^{-5}[{\rm eV}/{\rm K}]$.
@@ VI.1 C2H6, C2H4, and C2H5 molecules
Figures 4 (a)-(c) show the energy levels of C2H6, C2H4, and C2H5 molecular models, respectively. To draw figures of each local RCP on the electronic interfaces $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$, at least HOMO (Highest Occupied Molecular Orbital) should be included in the energy window $\mu_{\rm lower}<\mu<\mu_{\rm upper}$. Therefore, the energy windows were chosen to be -5 eV to 0 eV for C2H6, -3 eV to 0 eV for C2H4, and -2 eV to 0 eV for C2H5. The corresponding $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ are shown in Figs. 4 (d)-(f).
In Fig. 4 (d), although the H atoms in C2H6 are visualized as green parts, overall $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ is small. In contrast, the C2H4 (Fig. 4 (e)) has the high $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ (red regions) originating from a C-C double bond. This result is consistent with the fact that the double bonding region is a highly electron-donating region, where electrons are more likely to be supplied to outer regions. Figure 4 (f) shows a molecule model whose structure is generated as one H atom is deleted from C2H6 without geometry optimization. The red region at the top right in the C2H5 corresponds to the high $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ arising from a dangling bond that was formerly bonded to the deleted H atom. This fact is consistent with the HOMO distribution depicted in Fig. 4 (g).
@@
VI.2 Benzene, naphthalene, and pentacene molecules
The energy windows employed in the calculation of the RCPs $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ for benzene (C6H6), naphthalene (C10H8), and pentacene (C22H14) are shown in Figs. 5 (a)–(c), and the corresponding RCP distributions are presented in Figs. 5 (d)–(g). For pentacene, RCP $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ in the energy range between -2 eV and 0 eV is shown in Fig. 5 (e), and one in the energy range between 0 eV and 2 eV is shown in Fig. 5 (f), separately.
Benzene is known to have delocalized electrons due to the $\pi$ bonds created by the six-membered ring. Correspondingly, $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ is high throughout the benzene ring as shown in Fig. 5 (d). On the other hand, in the case of naphthalene shown in Fig. 5 (e), the high $\mu_{R}^{\tau}(\vec{r}_{\rm surf})$ due to the double bond is seen at four corners at the upper and lower ends. As easily expected, the RCP $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ originating from the Lowest Unoccupied Molecular Orbital (LUMO) contribution shown in Fig. 5 (f) is totally different from that originating from the HOMO contribution shown in Fig. 5 (e). Therefore, it is necessary to determine the appropriate energy window depending on whether the target substance is a donor or acceptor of electrons. It is clear from the $\mu_{R}^{\tau}(\vec{r})$ distribution shown in Fig. 5 (g) that pentacene has a bonding state between the outer central carbons similar to that of benzene rings, and four double bonding sites similar to those of naphthalene at the upper and lower ends, which is also consistent with the bond length argument. When molecules interact with another inactive material surface such as an AFM probe with CO molecules attached, high $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ becomes a measure of the Pauli repulsive force originating from antibonding. Figure 5 (g) is consistent with the topology image of non-contact AFM experiments using a CO probe [22, 45].
VI.3 Benzene derivatives
The energy levels, Hartree potentials at material surfaces, and the RCP $\mu_{R}^{\tau}(\vec{r}_{\rm{surf}})$ of phenol (C6H5OH) and nitrobenzene (C6H5NO2) molecules are shown in Fig. 6. It is known that substituted benzenes exhibit distinct regioselectivities in electrophilic aromatic substitution reactions. Phenol acts as an ortho/para-directing group due to the electron-donating effect of the hydroxyl substituent, whereas nitrobenzene is meta-directing as a result of the electron-withdrawing nature of the nitro group [46]. The electrostatic potential (Hartee potential), which is caused by intramolecular polarization, is a measure of Coulomb interaction with charged particles. Although the Hartree potential is sometimes used for discussions of electron-donating ability for ionic materials, it is not effective for regioselectivities in selectrophilic aromatic substitution reactions as shown in Figs. 6 (c) and (d). On the other hand, the local RCPs indicate the electron-donating ability at the surface. The distribution of local RCPs clearly visualizes the ortho/para-directing of the phenol and the meta-directing of the nitrobenzene. Here, we note that the energy windows were selected to include multiple orbitals, as it is known that the HOMO alone cannot account for the meta-directing behavior of nitrobenzene [47].